a right triangle gas an angle whose measure is 35. what is the measurment of the third angle?
Answer by ntnk(54)

There are two very important facts that you must know to do these types of questions.
1. In ANY triangle, the measurements of the three angles must add up to 180.
2. In a RIGHT triangle, one of the angles measures 90 degrees.
We have a right triangle with an angle measure of 35. The fact that the triangle is right tells us that one of the angles is 90. We need to find the measure of the third angle, which we will call x. Since the measures of the angles must add up to 180, we have
35 + 90 + x = 180
and we must solve for x. Simplifying a bit on the left hand side gives
125 + x = 180
Now subtract 125 from both sides to get
x = 180 - 125 = 55
so the third angle is 55 degrees. I hope this helps!